Fermi National Laboratory
 
Operation Summary for February 15, 2012 - February 17, 2012
 

Wednesday February 15

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, Operators halted NuMI beam at 8:57 AM for an access into their absorber room by Jasmin experimenters. A Water Group person reported at 9:30 AM that he was changing DI Bottle filters at MI-10. NuMI resumed taking beam at 9:46 AM. Operations searched and secured the MTest area around 11:35 AM. Operations established beam to the Meson MTest area for the EDIT School at 11:55 AM. At 1:30 PM, a Booster RF station (BRF11) tripped off for the first of three times in 45 minutes; after the third trip a RF person investigated. Booster and MI personnel began a joint phase-lock study at 2:35 PM.

During the evening shift, the joint Booster and MI studies ended at 4 PM. Operations disabled beam to the EDIT School at 5:52 PM. Operations established beam to the FTBF experiment T-979 at 6:11 PM. Operators reported around 7 PM that the MI tunnel temperature near a MI quadrupole magnet (Q305) alarmed high. Operations disabled beam to T-979 at 11:16 PM. At 11:49 PM, Operators prepared MI for an access to investigate a high water sump alarm in MI-40.

Thursday February 16

The midnight shift began with beam off to NuMI and MiniBooNE to investigate a MI sump pump alarm.

During the midnight shift, at 12:07 AM, Operators reported that the MI-40 sump was overflowing and Operators could not get the sump on; FESS personnel will investigate. At 12:51 AM, LRF5 tripped off; it reset. At 1:39 AM, FESS personnel reported that the MI-40 pump had failed and needed to be replaced; they will install a portable pump; the MI beam is off for this work. LRF4 tripped off at 5:53 AM; it reset. At 7:44 AM, Neutron Therapy Facility (NTF) personnel prepared to treat one patient later in the morning.

During the day shift, Water Group personnel reported at 8:02 AM that they were making up water for the MI-60 magnet LCW system. A Booster person began a 90-minute study at 8:04 AM. Operators found three MI vacuum pumps off at 9:27 AM; they all reset. The sump pump work ended and MI resumed sending beam at 2:10 PM. Operations established beam to the MTest EDIT School at 2:30 PM. Operators searched and secured the Neutrino Area G2 enclosure at 2:41 PM and tried to turn on power supplies, but only two of the eight supplies would run up; Operators contacted power supply personnel.

During the evening shift, Operations disabled beam to the MTest EDIT School at 5:56 PM. At 6:25 PM, a Linac Klystron RF station (KRF1) tripped off; it reset. At 7:14 PM, the NuMI beam permit tripped off due to an incomplete RF gap indication on MIRF18; everything reset. BRF15 and 16 tripped off at 7:30 PM; they reset. BRF16 tripped off again at 7:44 PM; it reset. Operations established beam to T-979 at 8:35 PM. LRF1 tripped off at 10:05 PM; it reset.

Friday February 17

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, Operations disabled beam to T-979 at 12:25 AM. Operators reported at 1:37 AM that they found a breaker in the off position in the G2 service building, which was holding off the power supplies; Operators got all the power supplies on but one, S:V424, which was being held off on a ground fault indication. Operators completed a 750 KeV line tune at 2:10 AM. Operators completed a minor tune change in MI at 3:32 AM. At 4:19 AM, the MI LCW system went into recirculation mode after column #1 at CUB went down; MI can stay in recirculation mode until the morning. At 6:36 AM, Operators prepared the NM2 enclosure for access. The NuMI beam permit tripped off at 6:53 AM; it reset. Operators talked with a Water Group person about column #1 at 7:23 AM; he said that column #2 was down for repairs and it would be a while before they could get to regenerate column #1, but the system should be fine in recirculation mode until Sunday – the good news was that in this configuration the MI LCW leak of about a hundred gallons a day went away. At 7:54 AM, NTF personnel prepared to treat three patients throughout the day.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo wants the complex to operate with as few beam disruptions as possible for the EDIT School today and for the other experiments throughout the weekend.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/17/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory